can you please help me with this
180-x=4(90-x)-15

Found 2 solutions by Nate, checkley71:
Answer by Nate(3500)

180-x=4(90-x)-15
180-x=360-4x-15
3x=165
x=55

Answer by checkley71(8403)

180-x=4(90-x)-15 or 180-x=360-4x-15 or 3x=360-15-180 or 3x=165 or x=165/3 or
x=55
proof 180-55=4(90-55)-15 or 125=360-220-15 or 125=360-235 or 125=125